# Sound velocity peak induced by the chiral partner in dense two-color QCD

### v1

@@ III.2 Thermodynamic quantities in LSM
The pressure within the LSM is straightforwardly evaluated from the potential (27). In terms of the physical quantities, $m_{\pi}$, $m_{\sigma}$ and $f_{\pi}$ provided in Eqs. (31), (32) and (33), it reads
@@ III.3 Sound velocity in LSM
We note that, the smaller value of $m_{\sigma}$ it is taken, the more rapidly the sound velocity converges on the conformal limit.
@@
by explicitly extracting Eq. (21). After this decomposition, $\delta c_{s}^{2}<0$ for any $\bar{\mu}$ $(\geq 1)$ is easily proven analytically while, of course, $\left(c_{s}^{\rm ChPT}\right)^{2}>0$. Hence, there appears a competition between $\left(c_{s}^{\rm ChPT}\right)^{2}>0$ and $\delta c_{s}^{2}<0$, i.e., between the ChPT contributions from the NG bosons and the correction provided from the chiral-partner dynamics. As a consequence, the emergence of non-monotonic behaviors in the sound velocity, including the peak structure, is expected.
@@ IV.1 Sound velocity
Depicted in Fig. 1 is the resultant $\mu_{q}$ dependences of the sound velocities evaluated in the LSM with several values of $m_{\sigma}$. The solid black curve corresponds to the ChPT result which is also reproduced within the LSM framework by taking $m_{\sigma}\to\infty$. The peak structure for certain choices of $m_{\sigma}$ is indeed observed, as the recent lattice data implies [30]. The figure also indicate that, in the vicinity of $\mu_{q}=\mu_{q}^{\rm cr}$, the results are less sensitive to the value of $m_{\sigma}$. In fact, from Eq. (48), the sound velocity for $\mu_{q}\approx\mu_{q}^{\rm cr}$ is simply approximated by
@@
which does not exhibit any $\delta\bar{m}_{\sigma-\pi}^{2}$ dependences. This linear line is shown by the dashed line in Fig. 1. On the other hand, when looking at the high-density region such as $\mu_{q}=3m_{\pi}\gg\mu_{q}^{\rm cr}$, the value of $c_{s}^{2}$ changes from $c_{s}^{2}=1$ toward $c_{s}^{2}=1/3$ in association with reduction of the $\sigma$ meson mass, indicating a significant dependence on $m_{\sigma}$. In fact, for finite $m_{\sigma}$, one can show that the sound velocity in high dense regime is expanded as
@@
This equation possesses a single solution of positive $\bar{\mu}_{p}$ only when $\delta\bar{m}^{2}_{\sigma-\pi}-8>0$ is satisfied. In other words, within the present LSM, we can obtain only one peak in the sound velocity and this peak is derived only when
@@
As the $\sigma$ meson mass becomes heavier, the suppression of the sound velocity from the correction $\delta c_{s}^{2}$ gets weakened. In order to gain insights into a relationship between this weakening and the peak properties, we depict the peak value $c_{s}^{2}(\bar{\mu}_{p})$ and the peak position $\bar{\mu}_{p}$ as functions of $m_{\sigma}$, in panels (a) and (b) of Fig. 3, respectively. The panel (a) of Fig. 3 shows that as $m_{\sigma}$ becomes larger from the critical value $m_{\sigma}=\sqrt{3}m_{\pi}$, the peak value is enhanced from $c_{s}^{2}=1/3$ toward $c_{s}^{2}=1$ monotonically. From the panel (b), one can see that there exists a lower limit of the peak position: $\bar{\mu}_{p}=2$, realized when $m_{\sigma}=3m_{\pi}$. In other words, the peak is generated only for $\mu_{q}>2\mu_{q}^{\rm cr}$ within the present LSM approach. Besides, the doubling of the peak position explained below Eq. 55 is clearly seen in this panel. Despite such a doubling, its peak value is distinct from the panel (a) of Fig. 3. Therefore, by combining both the informations, the peak value and the peak position, one can identify the vale of $m_{\sigma}$ exclusively.
@@ IV.2 Trace anomaly
$=$ $f_{\pi}^{2}m_{\pi}^{2}\left[\left\{\frac{16}{\delta\bar{m}^{2}_{\sigma-\pi}}\right\}(\bar{\mu}^{2}-1)\right],$
@@
From this expression, we find that in the high-density regions the sign of trace anomaly is related to the $\sigma$ meson mass. When $m_{\sigma}=\sqrt{3}m_{\pi}$ the first term of Eq. (60) vanishes, leading to $\bar{\Theta}_{\rm LSM}\to 6$ that is positive. For smaller $\sigma$ meson mass, $m_{\sigma}<\sqrt{3}m_{\pi}$, $\bar{\Theta}_{\rm LSM}$ positively diverges, meanwhile, when the $\sigma$ meson mass satisfies $m_{\sigma}>\sqrt{3}m_{\pi}$, the asymptotic sign of $\bar{\Theta}_{\rm LSM}$ turns out to be negative. It is noteworthy that the critical value for the sign of the trace anomaly $m_{\sigma}=\sqrt{3}m_{\pi}$ is identical to the one which discriminates the emergence of the sound velocity peak, as delineated in Sec. IV.1. Those remarkable correlations are summarized in Table 1.
$\sigma$ meson mass the peak of $c_{s}^{2}$ the asymptotic sign of $\bar{\Theta}$
$m_{\sigma}>\sqrt{3}m_{\pi}$ $\checkmark$ negative
@@ V Summary and discussion
In this paper, we have investigated the sound velocity at finite quark chemical potential and zero temperature in QC2D with two quark flavors, based on an effective model described by hadronic degrees of freedom. In particular, we have employed the linear sigma model, LSM, invented in Ref. [45] that is capable of describing not only the NG bosons associated with the chiral symmetry breaking but also their chiral partners as higher excitations, to go beyond the leading-order ChPT framework [6, 7].
Utilizing a mean-field approximation, we have analytically found that the thermodynamic quantities evaluated within the LSM are expressed by sums of the ChPT results and the corrections. Notably, the latter corrections are characterized by the inverse of a mass difference between the chiral partners, the $\sigma$ meson and pion. Hence, in a heavy mass limit of the $\sigma$ meson, the corrections vanish and the thermodynamic quantities in our LSM are reduced to the ChPT results, as naively expected. For this reason, in the QC2D thermodynamic aspects the LSM is regarded as a reasonable extension of the ChPT motivated by chiral symmetry. Therefore, we conclude that the LSM evaluation of the QC2D thermodynamics can serve as a new benchmark of hadron model approaches to compare with lattice observations.
@@
Our present investigation has been mostly devoted to examining the sound velocity in the hadronic level, utilizing the LSM treating spin-0 hadrons. As the chemical potential is increased, it is expected that further excitations such as spin-$1$ hadrons would also start to contribute to the sound velocity [49], and finally the quark degrees of freedom appear. However, how those other excited hadrons have influences on the sound velocity and how the quark-gluon dynamics begins to manifest themselves in the baryonic matter remain unclear [34, 31, 32, 33, 37, 50, 51], owing to the current limited lattice data in cold dense QC2D. Moreover, the $\sigma$ meson spectrum may be contaminated by a $0^{+}$ glueball through their mixings. Hence, in order to gain insights into those contributions, we expect more lattice results on not only the sound velocity but also the hadron spectrum in the future. Despite those uncertainties, our main harvest, the reproduction of the sound velocity peak in cold dense QC2D by extending the NG boson dynamics to include the chiral partners, would already have great significance.
@@ Acknowledgment
This work of M.K. is supported in part by the National Natural Science Foundation of China (NSFC) Grant Nos: 12235016, and the Strategic Priority Research Program of Chinese Academy of Sciences under Grant No XDB34030000. D.S. was supported by the RIKEN special postdoctoral researcher program and by the Japan Society for the Promotion of Science (JSPS) KAKENHI Grants No. 23K03377. The authors thank Naoki Yamamoto for fruitful discussions and comments, and Etsuko Itou and Kei Iida for useful information on the lattice data of the sound velocity and trace anomaly.

### v2

@@ III.2 Thermodynamic quantities in LSM
The pressure within the LSM is straightforwardly evaluated from the potential (27). In terms of the physical quantities, $m_{\pi}$, $m_{\sigma}$ and $f_{\pi}$ provided in Eqs. (31), (32), and (33), it reads
@@ III.3 Sound velocity in LSM
We note that the smaller value of $m_{\sigma}$ taken the more rapidly the sound velocity converges on the conformal limit.
@@
by explicitly extracting Eq. (21). After this decomposition, $\delta c_{s}^{2}<0$ for any $\bar{\mu}$ $(\geq 1)$ is easily proven analytically while, of course, $\left(c_{s}^{\rm ChPT}\right)^{2}>0$. Hence, there appears a competition between $\left(c_{s}^{\rm ChPT}\right)^{2}>0$ and $\delta c_{s}^{2}<0$, i.e., between the ChPT contributions from the NG bosons and the correction provided from the chiral-partner dynamics. As a consequence, the emergence of nonmonotonic behaviors in the sound velocity, including the peak structure, is expected.
@@ IV.1 Sound velocity
Depicted in Fig. 1 is the resultant $\mu_{q}$ dependencies of the sound velocities evaluated in the LSM with several values of $m_{\sigma}$. The solid black curve corresponds to the ChPT result which is also reproduced within the LSM framework by taking $m_{\sigma}\to\infty$. The peak structure for certain choices of $m_{\sigma}$ is indeed observed, as the recent lattice data implies [32]. The figure also indicate that, in the vicinity of $\mu_{q}=\mu_{q}^{\rm cr}$, the results are less sensitive to the value of $m_{\sigma}$. In fact, from Eq. (48), the sound velocity for $\mu_{q}\approx\mu_{q}^{\rm cr}$ is simply approximated by
@@
which does not exhibit any $\delta\bar{m}_{\sigma-\pi}^{2}$ dependencies. This linear line is shown by the dashed line in Fig. 1. On the other hand, when looking at the high-density region such as $\mu_{q}=3m_{\pi}\gg\mu_{q}^{\rm cr}$, the value of $c_{s}^{2}$ changes from $c_{s}^{2}=1$ toward $c_{s}^{2}=1/3$ in association with reduction of the $\sigma$ meson mass, indicating a significant dependence on $m_{\sigma}$. In fact, for finite $m_{\sigma}$, one can show that the sound velocity in high dense regime is expanded as
@@
This equation possesses a single solution of positive $\bar{\mu}_{p}^{2}$ only when $\delta\bar{m}^{2}_{\sigma-\pi}-8>0$ is satisfied. In other words, within the present LSM, we can obtain only one peak in the sound velocity and this peak is derived only when
@@
As the $\sigma$ meson mass becomes heavier, the suppression of the sound velocity from the correction $\delta c_{s}^{2}$ gets weakened. In order to gain insights into a relationship between this weakening and the peak properties, we depict the peak value $c_{s}^{2}(\bar{\mu}_{p})$ and the peak position $\bar{\mu}_{p}$ as functions of $m_{\sigma}$, in panels (a) and (b) of Fig. 3, respectively. Panel (a) of Fig. 3 shows that as $m_{\sigma}$ becomes larger from the critical value $m_{\sigma}=\sqrt{3}m_{\pi}$, the peak value is enhanced from $c_{s}^{2}=1/3$ toward $c_{s}^{2}=1$ monotonically. From panel (b), one can see that there exists a lower limit of the peak position, $\bar{\mu}_{p}=2$, realized when $m_{\sigma}=3m_{\pi}$. In other words, the peak is generated only for $\mu_{q}>2\mu_{q}^{\rm cr}$ within the present LSM approach. Besides, the doubling of the peak position explained below Eq. (55) is clearly seen in this panel. Despite such a doubling, its peak value is distinct from panel (a) of Fig. 3. Therefore, by combining both the information, the peak value and the peak position, one can identify the value of $m_{\sigma}$ exclusively.
@@ IV.2 Trace anomaly
$=$ $f_{\pi}^{2}m_{\pi}^{2}\left[\frac{16}{\delta\bar{m}^{2}_{\sigma-\pi}}(\bar{\mu}^{2}-1)\right],$
@@
From this expression, we find that in the high-density regions the sign of trace anomaly is related to the $\sigma$ meson mass. When $m_{\sigma}=\sqrt{3}m_{\pi}$ the first term of Eq. (60) vanishes, leading to $\bar{\Theta}_{\rm LSM}\to 6$ that is positive. For smaller $\sigma$ meson mass, $m_{\sigma}<\sqrt{3}m_{\pi}$, $\bar{\Theta}_{\rm LSM}$ positively diverges; meanwhile, when the $\sigma$ meson mass satisfies $m_{\sigma}>\sqrt{3}m_{\pi}$, the asymptotic sign of $\bar{\Theta}_{\rm LSM}$ turns out to be negative. It is noteworthy that the critical value for the sign of the trace anomaly $m_{\sigma}=\sqrt{3}m_{\pi}$ is identical to the one which discriminates the emergence of the sound velocity peak, as delineated in Sec. IV.1. Those remarkable correlations are summarized in Table 1.
$\sigma$ Meson mass Peak of $c_{s}^{2}$ Asymptotic sign of $\bar{\Theta}$
$m_{\sigma}>\sqrt{3}m_{\pi}$checkmark Negative
@@ V Summary and discussion
In this paper, we have investigated the sound velocity at the finite quark chemical potential and zero temperature in QC2D with two quark flavors, based on an effective model described by hadronic degrees of freedom. In particular, we have employed the linear sigma model, LSM, invented in Ref. [55] that is capable of describing not only the NG bosons associated with the chiral symmetry breaking but also their chiral partners as higher excitations, to go beyond the leading-order ChPT framework [6, 7].
Utilizing a mean-field approximation, we have analytically found that the thermodynamic quantities evaluated within the LSM are expressed by sums of the ChPT results and the corrections. Notably, the latter corrections are characterized by the inverse of a mass difference between the chiral partners, the $\sigma$ meson and pion. Hence, in a heavy mass limit of the $\sigma$ meson, the corrections vanish and the thermodynamic quantities in our LSM are reduced to the ChPT results, as naively expected. For this reason, in the QC2D thermodynamic aspects, the LSM is regarded as a reasonable extension of the ChPT motivated by chiral symmetry. Therefore, we conclude that the LSM evaluation of the QC2D thermodynamics can serve as a new benchmark of hadron model approaches to compare with lattice observations.
@@
Our findings are evaluated within the mean-field approximation. To confirm the validity of our evaluations at the quantum level, it is worth going beyond the mean-field approximation in the LSM. Indeed, the one-loop calculation of the NG-boson fluctuations has already been taken into account in the ChPT analysis [59, 60, 61]. Given this fact, the inclusion of the chiral partner contribution in the one-loop calculation would be straightforward. More explicit analysis of such loop corrections in the LSM is to be pursued elsewhere.
Our present investigation has been mostly devoted to examining the sound velocity in the hadronic level, utilizing the LSM treating spin-00 hadrons. As the chemical potential is increased, it is expected that further excitations such as spin-$1$ hadrons would also start to contribute to the sound velocity [62], and finally the quark degrees of freedom appear. However, how those other excited hadrons have influences on the sound velocity and how the quark-gluon dynamics begins to manifest themselves in the baryonic matter remain unclear [41, 38, 39, 40, 44, 47, 48], owing to the current limited lattice data in cold dense QC2D. Moreover, the $\sigma$ meson spectrum may be contaminated by a $0^{+}$ glueball through their mixings. Hence, in order to gain insights into those contributions, we expect more lattice results on not only the sound velocity but also the hadron spectrum in the future. Despite those uncertainties, our main harvest, the reproduction of the sound velocity peak in cold dense QC2D by extending the NG boson dynamics to include the chiral partners, would already have great significance.
@@ ACKNOWLEDGMENT
This work of M.K. is supported in part by the National Natural Science Foundation of China (NSFC) Grant No. 12235016, and the Strategic Priority Research Program of Chinese Academy of Sciences under Grant No XDB34030000. D.S. was supported by the RIKEN special postdoctoral researcher program and by the Japan Society for the Promotion of Science (JSPS) KAKENHI Grants No. 23K03377 and 23H05439. The authors thank Naoki Yamamoto for fruitful discussions and comments and Etsuko Itou and Kei Iida for useful information on the lattice data of the sound velocity and trace anomaly.
Appendix A Spontaneous chiral symmetry breaking in low-energy QC2D and LSM
In this appendix, we derive the connection between the chiral condensate and pion decay constant in Eq. (33).
We start from the underlying Lagrangian of the two-flavor QC2D,
${\cal L}_{\rm QC_{2}D}=\bar{\psi}(i\gamma^{\mu}D_{\mu}-m_{l})\psi-\frac{1}{4}G_{\mu\nu}^{a}G^{\mu\nu,a},$ (61)
where $\psi=(u,d)^{T}$ denotes the two-flavor quark doublet, and $D_{\mu}\psi=(\partial_{\mu}+igA_{\mu}^{a}T_{c}^{a})\psi$ is the covariant derivative incorporating interactions with a gluon field $A_{\mu}^{a}$. The $2\times 2$ matrix $T_{c}^{a}=\tau_{c}^{a}/2$ is the generator of $SU(2)_{c}$ color group with $\tau_{c}^{a}$ being the Pauli matrix; $g$ and $m_{l}$ are the QC2D coupling constant and an isospin symmetric mass of current quarks, $m_{u}=m_{d}\equiv m_{l}$. For the sake of convenience in the following discussion, we take the Weyl representation of the Dirac matrices,
$\gamma^{\mu}$ $=$ $\begin{pmatrix}0&\sigma^{\mu}\\ \bar{\sigma}^{\mu}&0\end{pmatrix},\;\;\;\gamma_{5}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},$ (62)
where $\sigma^{\mu}=({\bm{1}},\sigma^{i})$ and $\bar{\sigma}^{\mu}=({\bm{1}},-\sigma^{i})$ with the Pauli matrix $\sigma^{i}$ in spinor space. By using the Weyl representation, the quark field, which belongs to the fundamental representation for $SU(2)_{c}$, is decomposed into the left- and right-handed components,
$\psi=(\psi_{R},\psi_{L})^{T}.$ (63)
From the pseudoreal viewpoint of $SU(2)$ groups, it is convenient to introduce the following “conjugate fields”,
$\tilde{\psi}_{L,R}$ $=$ $\sigma_{2}\tau_{c}^{2}\psi^{*}_{L,R}$ (64)
$=$ $(\tilde{u}_{L,R},\tilde{d}_{L,R})^{T}.$
Note that this pseudoreal representation satisfies a property similar to the Grassmann property for fundamental quarks $\psi_{L(R)}^{T}\psi_{L(R)}=0$,
$\tilde{\psi}_{L(R)}^{\dagger}\psi_{L(R)}=0.$ (65)
By combining the fundamental and pseudoreal representations, one can define the following field,
$\Psi=(\psi_{R},\tilde{\psi}_{L})^{T}=(u_{R},d_{R},\tilde{u}_{L},\tilde{d}_{L})^{T},$ (66)
where these four components are in the flavor space. Then, the two-flavor QC2D Lagrangian can be expressed as
${\cal L}_{\rm QC_{2}D}=\Psi^{\dagger}\sigma^{\mu}\left(i\partial_{\mu}-gA_{\mu}^{a}T_{c}^{a}\right)\Psi-\frac{1}{4}G_{\mu\nu}^{a}G^{\mu\nu,a}$
$\;\;\;\;\;\;\;\;\;\;\;\;-\frac{m_{l}}{2}\left(\Psi^{T}\sigma^{2}\tau_{c}^{2}E^{T}\Psi+({\rm H.c.})\right).$ (67)
At the massless limit of the current quark masses, this Lagrangian is invariant under the $SU(4)$ transformation,
$\Psi\to g\Psi\;\;\mbox{with $g\in SU(4)$}.$ (68)
By using the $\Psi$ field, the quark condensate can be described as
$\langle\bar{\psi}\psi\rangle=\langle\left(\frac{1}{2}\Psi^{T}\sigma^{2}\tau_{c}^{2}E^{T}\Psi+({\rm H.c.})\right)\rangle.$ (69)
This quark condensate constructed by $\Psi$ is not invariant under the $SU(4)$ transformation but invariant under the $Sp(4)$ transformation,
ΨhΨwith $\Psi\to h\Psi\;\;\mbox{with $h\in Sp(4)$}.$ (70)
Namely, when the quark condensate gets a finite value, the $SU(4)$ symmetry is spontaneously broken into the $Sp(4)$ symmetry in the vacuum of the two-flavor QC2D. Consequently, five NG bosons show up, which is associated with the number of the broken generators of $SU(4)/Sp(4)$, $X^{i}$ ($i=1\sim 5$). Indeed, a current of the broken symmetry is coupled to the NG bosons.
Under the broken symmetry transformation with the rotation angle $\theta^{i}$, the quark field $\Psi$ is transformed as
$\Psi\to\exp[-i\theta^{i}X^{i}]\Psi.$ (71)
Then, one can obtain the broken current,
$j_{5}^{i\mu}$ $=$ $\Psi^{\dagger}\sigma_{\mu}X^{i}\Psi.$ (72)
Note that this current can be rewritten by the original field $\psi$,
$j_{5}^{i\mu}=$
$\begin{cases}\frac{1}{2\sqrt{2}}\bar{\psi}\gamma^{\mu}\gamma_{5}\tau^{i}_{f}\psi\;\;(\mbox{for }i=1,2,3)\\ \frac{i}{4\sqrt{2}}\psi^{T}C\gamma_{5}\tau^{2}_{c}(D_{f}^{i})^{\dagger}(\gamma^{\mu})^{\dagger}\psi+({\rm H.c.})\;\;(\mbox{for }i=4,5)\end{cases}$
where $C=i\gamma^{2}\gamma^{0}$ denotes the charge-conjugation operator. As a consequence of the spontaneous chiral symmetry breaking, there exists an overlap amplitude between the broken current $j^{i\mu}_{5}$ and the NG bosons $\mathscr{P}^{i}$ in the low-energy regime of QC2D,
$\langle 0|j^{i\mu}_{5}(x)|\mathscr{P}^{j}(p)\rangle=-if_{\pi}p^{\mu}e^{-ip\cdot x}\delta^{ij}\;\;\;(\mbox{$i,j=1\sim 5$}),$
where $f_{\pi}$ denotes the pion decay constant.
In the LSM, we can also define the broken current at the vacuum ($\mu_{q}=0$). By following the underlying transformation in Eq. (71), $\Sigma$ also transforms under the broken symmetry of $SU(4)$ as
$\Sigma\to\exp[-i\theta^{i}X^{i}]\Sigma\exp[-i\theta^{i}(X^{i})^{T}].$ (75)
From Eq. (24), the broken current of the LSM reads
$j_{5}^{i\mu}=2(\partial^{\mu}\mathscr{P}^{j}\mathscr{S}^{k}-\partial^{\mu}\mathscr{S}^{j}\mathscr{P}^{k}){\rm tr}\Bigl{[}\{X^{j},X^{k}\}X^{i}\Bigl{]}.$ (76)
In the hadronic phase where the vacuum undergoes the spontaneous chiral symmetry breaking, $\sigma_{0}=\langle\mathscr{S}^{0}\rangle\ \neq 0$ and $\Delta=\langle\mathscr{P}^{5}\rangle=0$, the broken current takes the form of
$j_{5}^{i\mu}$ $=$ $\frac{1}{\sqrt{2}}\sigma_{0}\partial^{\mu}\mathscr{P}^{i}+\cdots.$ (77)
Here we have picked up a term relevant to NG bosons. Thus, within the LSM, the overlap amplitude between the broken current $j_{5}^{i\mu}$ and the NG bosons $\mathscr{P}^{i}$ is evaluated as
$\langle 0|j^{i\mu}_{5}(x)|\mathscr{P}^{j}(p)\rangle=-i\frac{1}{\sqrt{2}}\sigma_{0}p^{\mu}e^{-ip\cdot x}\delta^{ij}.$ (78)
Matching to the underlying definition in Eq. (74), the chiral condensate can be connected to the pion decay constant as
$\sigma_{0}=\sqrt{2}f_{\pi}.$ (79)
Note that the factor $\sqrt{2}$ deviates from the well-known case of the three-color QCD with two flavors since the chiral $SU(2)$ symmetry is extended to the $SU(4)$ symmetry due to the pseudoreality of QC2D.